# HiFi-Stream: Streaming Speech Enhancement with Generative Adversarial Networks

### v1

@@ V-A Offline Comparison (Full Audio)
All the modifications were trained with the same computational resource, the results are depicted in the Table I. HiFi++(reprod) has lower metrics than presented in the repository due to the less training time invested. We compare the modified models to it to have the same training conditions. HiFi w/o spec acts as a starting point. The usage of 2D convolutions together with optimization of number of channels (HiFi-2dMRF) allowed a drastic reduction in parameters of HiFi Block resulting in almost 3x parameter reduction in comparison to HiFi+w/o spec. It presents the direction of smaller number of parameters but in cost of computational complexity.
The usage of FMS allows not only the reduced complexity and parameters but also better performance in terms of quality. HiFi-fms is essentially HiFi++ w/o spec with FMSBlock in place of MaskNet and HiFi-fms+2dMRF additionally uses 2D convolutions in HiFi MRF Blocks. The two models have their effect in comparison to HiFi++(reprod.) and HiFi-2dMRF depicting the effect of FMS Blocks. As we can see from the tables, the modifications with FMS despite the reduced complexity demonstrate relatively the same metrics or better than HiFi w/o spec and HiFi++(reprod.).
Overall, HiFi-Stream2D model can be chosen as a small-size model and HiFi-Stream is the better option in terms of computational complexity, according to the MACs, see Fig.3.
We also compared our solutions to the available baselines including some recent GAN-based light-weight solutions. According to Table I, HiFi-Stream models demonstrate similar performance metrics with a slight drop in comparison to existing low-parameter baselines. In particular, the model reaches the same results as the lightest FFC-AE-V0 with almost the same number of parameters but 2x reduced computational complexity.
@@ V-B Streaming Comparison
We compare the stream-aware models in the streaming conditions (see Table II) without additional streaming training. It is notable that almost in all cases the quality declines in all metrics which is the effect of limited audio input and context. In comparison to published baselines HiFi++ models with FMS blocks demonstrate their robustness in this regard showing slightly worse performance. It may be the case that due to the filter-wise scaling in FMS mechanism it better exploits short-time correlations while other models may require additional streaming finetuning to compensate for the lost global context.
@@ VI Conclusion
Our results show that HiFi++ model can be optimized to become another GAN candidate to be the most optimal trade-off between complexity and quality. Among considered GAN baselines the proposed solutions are the best in terms of computational complexity. The achieved results suggest that the model is robust to limited context conditions and therefore can be used in SE streaming applications.

### v2

@@ V-A Offline Comparison (Full Audio)
The results of the offline evaluation on the test set are shown in Table I. The top part presents baseline models from the literature while the bottom part shows our own trained models.
Paper description differs from the checkpoint presented in HiFi++ repository in the number of convolution channels in the HiFi block. It has slightly different parameter count and complexity. Also the github checkpoint was trained for longer time and showed better metrics. We compare the modified models to our reproduction (named HiFi++(reprod)) to have the same training conditions for fair comparison but also display the original checkpoint results (denoted as HiFi++, top part).
HiFi-w/o-spec is HiFi++(reprod) without Spectral U-Net and other modifications are built on top of this model. From the results we observe that model HiFi-2dMRF with time-frequency 2D convolutions and reduced number of convolution channels showed drastic reduction in parameters of HiFi Block resulting in almost 3x parameter reduction in comparison to HiFi-w/o-spec.
The effect of FMS is presented by models HiFi-Stream and HiFi-Stream2D. The former is HiFi-w/o-spec model with FMS MaskNet in the end. It allowed not only the reduced complexity and parameters but also better quality. HiFi-Stream2D possesses additionally 2D convolutions in HiFi MRF blocks and it allowed even further parameter reduction.
As we can see from the Table I, despite the reduced complexity, the modifications with FMS (HiFi-Stream and HiFiStream2D) demonstrate relatively the same metrics or better than HiFi w/o spec and HiFi++(reprod.).
We compared our solutions to the available baselines including some recent GAN-based light-weight solutions. According to Table I, HiFi-Stream models demonstrate similar performance metrics in comparison to existing low-parameter baselines. In particular, the model reaches the same results as the lightest FFC-AE-V0 with almost the same number of parameters but two times reduced computational complexity.
@@ V-B Streaming Comparison
We compare the stream-aware models in the streaming conditions (see Table II) without additional streaming training. It can be noted that the quality declines compared to offline evaluation in all metrics which is the effect of limited audio input and context. In comparison to published baselines HiFi++ models with FMS blocks demonstrate their robustness in this regard showing slightly worse performance. It demonstrates that FMS MaskNet exploits the given context more efficiently than U-Net in the original Spectral MaskNet.
Overall, HiFi-Stream2D model can be chosen as a small-size model and HiFi-Stream is the better option in terms of computational complexity, according to the MACs. To illustrate it more, we put the results in Fig.3 showing the complexity metrics against perceptual quality given by COVL. Our models HiFi-Stream and HiFi-Stream2D clearly stand as better complexity-quality trade-offs than considered baselines.
@@ VI Conclusion
In our work we propose two modification of HiFi++ model serving as better complexity-quality trade-off. Our results indicate that the proposed methods are more robust in changing context that makes them better suited for streaming applications.